# Recent results from MAGIC

### v1

@@ 2.5 Cosmic-ray electrons and positrons
The cosmic-ray electrons and positrons (CREs) make up less than 1% of the total cosmic-ray flux, composed mainly of cosmic-ray protons. Due to the low mass of the electrons and positrons, they experience significant radiative losses and these losses increase quickly with increasing particle energy. As a result, the tortuous trajectory in the Galactic magnetic field of these TeV CREs are expected to reach a distance from the source of less the 1 kiloparsec. Therefore such a CREs measurement gives a good indication of existing nearby sources. MAGIC has used machine learning classification to distinguish CREs from protons and heavier nuclei. By analyzing 220 hours of MAGIC data, a CRE spectrum spanning an energy range from 500 GeV to 6 TeV has been measured [16] and aligns well with the results from other experiments, like H.E.S.S. [17], VERITAS [18], DAMPE [19, 20], AMS [21] and CALET [22].

### v2

@@ 2.5 Cosmic-ray electrons and positrons
The cosmic-ray electrons and positrons (CREs) make up less than 1% of the total cosmic-ray flux, composed mainly of cosmic-ray protons. Due to the low mass of the electrons and positrons, they experience significant radiative losses and these losses increase quickly with increasing particle energy. As a result, the tortuous trajectory in the Galactic magnetic field of these TeV CREs are expected to reach a distance from the source of less than 1 kiloparsec. Therefore such a CREs measurement gives a good indication of existing nearby sources. MAGIC has used machine learning classification to distinguish CREs from protons and heavier nuclei. By analyzing 220 hours of MAGIC data, a CRE spectrum spanning an energy range from 300 GeV to 6 TeV has been measured [16] and aligns well with the results from other experiments, like H.E.S.S. [17], VERITAS [18], DAMPE [19, 20], AMS [21] and CALET [22].